# Establishing Minimum Elements for Effective Vulnerability Management in AI Software

### v1

@@ Abstract
In the rapidly evolving field of artificial intelligence (AI), the identification, documentation, and mitigation of vulnerabilities are paramount to ensuring robust and secure systems. This paper discusses the minimum elements for AI vulnerability management and the establishment of an Artificial Intelligence Vulnerability Database (AIVD). It presents standardized formats and protocols for disclosing, analyzing, cataloging, and documenting AI vulnerabilities. It discusses how such an AI incident database must extend beyond the traditional scope of vulnerabilities by focusing on the unique aspects of AI systems. Additionally, this paper highlights challenges and gaps in AI Vulnerability Management, including the need for new severity scores, weakness enumeration systems, and comprehensive mitigation strategies specifically designed to address the multifaceted nature of AI vulnerabilities.
Index Terms:
AI Incidents, AIVD, AI Vulnerability Management, AI Bill of Materials, AIBOM
@@ I Introduction
As artificial intelligence (AI) technologies permeate various sectors including healthcare [1], finance [2], transportation [3], industry[4, 5], and security [6, 7, 8], their transformative impact is undeniable. However, this rapid integration introduces complex vulnerabilities that may compromise the systems it is intended to enhance. With AI becoming a foundational component across various sectors, there is an urgent need to systematically identify, evaluate, and mitigate AI-specific vulnerabilities. Therefore, Vulnerability Management in AI systems is a critical process that encompasses the identification, assessment, public disclosure, and remediation of security vulnerabilities within the systems and components. This practice is essential due to the unique and complex nature of AI software, which often includes data-driven decision-making processes that are opaque and challenging to secure.
@@
The establishment of a robust vulnerability management framework for AI systems underscores the need for centralized information sharing, where AI vulnerabilities can be reported, cataloged, and communicated with all stakeholders. This centralization ensures that information regarding new and emerging threats is quickly and widely disseminated, allowing for more rapid responses across the industry. This demands a global AI vulnerability database to serve as a central repository for information on identified vulnerabilities specific to AI systems, providing crucial support for organizations in protecting their AI technologies against security threats. While the National Vulnerability Database (NVD) [9], managed by the National Institute of Standards and Technology (NIST), offers a proven framework for cataloging software vulnerabilities, the entire vulnerability management ecosystem, is not designed to scale to support AI vulnerabilities and incidents.
@@
AI vulnerabilities differ significantly from traditional software vulnerabilities due to the adaptive and often data-driven nature of AI systems. Unlike conventional software, where vulnerabilities typically arise from coding errors or system misconfigurations, AI vulnerabilities often stem from the training data, model architecture, or the learning algorithms themselves. For instance, adversarial attacks[10], where malicious inputs are designed to fool AI models into making incorrect decisions, highlight vulnerabilities unique to AI systems. These vulnerabilities can be exploited to cause AI systems to behave unpredictably or to make decisions that benefit attackers. Understanding these vulnerabilities requires a deep dive into the underlying algorithms and data, making the process of identifying and mitigating these vulnerabilities complex and challenging.
Given the unique nature of AI vulnerabilities, there is a clear need for a specialized repository that can serve as a central hub for information on AI-specific vulnerabilities. The proposed Artificial Intelligence Vulnerability Database (AIVD) would function similarly to the NVD but would be specifically tailored to catalog and disseminate information about vulnerabilities unique to AI systems. In this paper, we discuss the minimum elements for AI vulnerability management and establishing an Artificial Intelligence Vulnerability Database. We introduce standardized formats and protocols for disclosing, analyzing, cataloging, and measuring AI vulnerabilities. We discuss how such AI incident database must extend beyond the traditional scope of the NVD by focusing on the unique aspects of AI systems. It needs to catalog issues such as biases in training data, flaws in algorithmic design, and vulnerabilities specific to AI deployment environments. Additionally, we highlight challenges and gaps in AI vulnerability management such as the need for new severity scores, weakness enumeration systems, and comprehensive mitigation strategies, specifically designed to address the multifaceted nature of AI vulnerabilities. These systems, dependent on data quality and algorithmic complexity, require adaptive approaches due to their dynamic, non-deterministic nature. SectionVI further discusses the gaps and challenges in AI vulnerability management.
@@
1 AI-CVE ID This is a unique identifier assigned to each recorded vulnerability within the AIVD to ensure easy tracking and referencing.
2 AI Model Details Specific details about the AI model or system affected (e.g., model type, version).
3 Weakness Type Identifies the type of vulnerability found in the AI system, providing details on the general category or the specific mechanism by which the system is compromised. Similar to CWE in NVD database, each vulnerability will be assigned with an AI-CWE.
4 Root Cause Explains the underlying factors contributing to the vulnerability. This root cause could be due to issues such as inherent vulnerabilities within the model, inadequate resiliency and robustness safeguards, and flaws in data.
@@
6 Severity Scores Adapted to reflect the unique impacts and severity considerations of AI vulnerabilities. Determining the severity score is crucial because it quantifies the severity of vulnerabilities, guiding prioritization and response strategies.
7 Affected Software & Products Applications susceptible to the vulnerability, as it allows organizations to precisely target their remediation efforts, manage risks effectively, and ensure compliance with security standards.
8 Exploitability It explains how a vulnerability can be exploited, focusing on the technical complexity of the attack, the level of privileges necessary, and the specific actions required by users to exploit the vulnerability.
@@ II Minimum Elements for Defining AI Vulnerabilities
This section describes the Minimum Elements (MEs) essential for capturing and describing AI vulnerabilities effectively. Each ME represents a critical piece of information that contributes to the comprehensive profiling and understanding of vulnerabilities within AI systems. These elements facilitate structured data collection, enabling systematic tracking, analysis, and mitigation of vulnerabilities, and equip stakeholders with necessary details to effectively address vulnerabilities from identification through resolution. Each ME is strategically defined to encapsulate crucial information regarding the vulnerabilities it describes. This approach not only standardizes vulnerability reporting across various platforms and industries but also ensures that all stakeholders—from developers and security professionals to regulatory bodies—have access to the same comprehensive data set. The structured data collection enabled by these elements allows organizations to systematically track vulnerabilities from initial identification to final resolution. This process includes detailed analysis stages where vulnerabilities are assessed for impact and exploitability, allowing for informed prioritization in mitigation efforts.
Furthermore, these elements serve as the backbone for creating databases and registries that stakeholders can reference to ensure consistency and accuracy in vulnerability management. By defining these minimum required data points, the AI vulnerability disclosure process becomes more efficient, transparent, and scalable, accommodating the dynamic nature of AI technologies and the complex ecosystems in which they operate. This standardized approach not only streamlines the mitigation process but also enhances the security posture of AI systems by providing clear and actionable information that can be rapidly deployed to counter potential threats. These minimum elements are described in the Table I. Some of these minimum elements require further advancements in foundational knowledge and the creation of new frameworks, and catalogs to describe AI vulnerabilities.
@@ II-A Weakness Type
Despite their sophisticated capabilities, AI systems are uniquely vulnerable to certain types of weaknesses that are seldom encountered in traditional software systems. These vulnerabilities stem from inherent characteristics of AI, such as heavy reliance on expansive datasets and complex decision-making processes that lack transparency. While resources like CWE exist to catalog vulnerabilities in traditional software systems, there is currently no equivalent resource specifically tailored for AI. Moreover, the CWE framework cannot fully capture the unique weaknesses of AI systems, such as those introduced by their reliance on data and algorithmic complexity. AI’s dependency on data quality and volume means that any corruption or bias in the data can significantly impact performance and decision accuracy. Furthermore, the black-box nature of many AI algorithms complicates the detection and understanding of vulnerabilities, making it challenging to develop robust AI systems.
Building on the need to address the unique vulnerabilities of AI systems, we propose the development of an AI-specific Common Weakness Enumeration (AI-CWE). AI-CWE will comprehensively catalog and enumerate AI-specific weaknesses, thereby facilitating a more structured approach to identifying these vulnerabilities. To structure the AI-CWE, we draw upon the AI/ML ATT&CK framework[11], ensuring that each identified weakness is linked to a specific deficiency in the system’s mitigation strategies. For instance, the four main classes of weaknesses are as follows:
@@
Each of these weaknesses will be further broken down into more detailed categories for analysis and mitigation. This categorization allows users to better understand the specific type of weakness within their AI system, enabling them to reason about the root cause of the weakness and employ more targeted and effective mitigation strategies. Tables II and III present two samples of AI-CWEs and the components to consider in describing them.
@@
TABLE II: Example 1- A potential AI-CWE entry: Inadequate Input Filtering
AI-CWE-ID AI-CWE-101
AI-CWE-Name Lack of Appropriate Privacy Safeguard
Description Privacy safeguards are deficient or absent, exposing the data to unauthorized access and manipulation due to inadequate encryption, anonymization, or access control configurations.
Examples The lack of appropriate privacy safeguards allows malicious actors to perform black-box attacks against the system to steal private information about the system
Severity Critical
Common Consequence Information Disclosure
Relationships None
Weakness Phase Inference, Training
Potential Mitigation Differential Privacy & Prediction Restriction
References https://ieeexplore.ieee.org/abstract/document/9745062, https://ieeexplore.ieee.org/document/10214061 [14], https://proceedings.neurips.cc/paper_files/paper/2023/hash/e9df36b21ff4ee211a8b71ee8b7e9f57-Abstrac.t-Conference.html [15]
TABLE III: Example 1- A potential AI-CWE entry: Lack of Privacy Safeguard
@@ II-B Impact
Describes the repercussions of a vulnerability within the AI system, detailing how the flaw affects the functionality, security, and operational reliability. It evaluates both direct and indirect effects, such as system performance degradation, potential exploitation risks, and broader implications for system users and associated services.
@@ II-C Severity Scores
The Common Vulnerability Scoring System (CVSS) is a widely used standard for assessing the severity of security vulnerabilities in software. However, in the context of AI systems, the unique characteristics of these technologies expose several gaps in the CVSS and other similar framework. Given the distinctive nature of AI vulnerabilities, a tailored approach to severity assessment is necessary. Below, we highlight the gaps in current methodologies like CVSS, discuss how to measure the severity of AI vulnerabilities, and propose viable directions.
There are several challenges in severity measurement of AI vulnerabilities.
@@ II-C1 CVSS Base Metrics and AI
  • Exploitability Metrics (Attack Vector, Attack Complexity, Privileges Required, User Interaction): These metrics provide a fundamental understanding of how a vulnerability can be exploited. However, AI systems may introduce additional complexities not captured by these metrics, such as the subtle ways data manipulation can influence AI behavior.
  • Impact Metrics (Confidentiality, Integrity, Availability): While these metrics address direct impacts of vulnerabilities, AI systems might suffer from indirect effects such as decision-making errors or biases introduced through data poisoning, which are not directly related to data loss or system downtime, but affect the system’s operational integrity.
@@ II-C2 Environmental Metrics and AI
These metrics allow adaptation to particular organizational infrastructures, which is crucial for AI systems as their deployment contexts can significantly influence vulnerability impacts. However, the dynamics of AI behavior, learning over time, and interaction with varying data streams might require even more dynamically adjustable metrics.
@@ II-C3 Supplemental Metrics and AI
Safety, Automatable, Recovery, and Value Density metrics are designed to provide additional context. For AI systems, the safety metric is particularly crucial as AI decisions might directly impact human safety. The automatable metric is especially important in the context of AI, considering the potential for automated attacks exploiting AI vulnerabilities across multiple systems.
@@ II-C4 Notable Gaps
Despite the comprehensive nature of CVSS, there are several notable gaps when applied to AI: (1) Dynamic Nature of AI: AI systems evolve, learning from new data, which might alter their vulnerability landscape over time. CVSS scores, being largely static once assigned, may not reflect the changing nature of AI vulnerabilities. (2) Indirect Impacts: AI vulnerabilities might not directly compromise data confidentiality, integrity, or availability but could degrade the quality of AI decisions, leading to erroneous outcomes. These impacts are not well captured by current CVSS metrics. (3) Data and Model Specific Attacks: Attacks on AI may involve data manipulation or model evasion techniques that do not have direct analogs in traditional software vulnerabilities, requiring new types of impact and exploitability metrics.
II-C5 Considerations
Adapting the CVSS to better fit the characteristics of AI vulnerabilities involves a deeper examination of how AI-specific risks can be systematically and dynamically assessed.
  • Dynamic Scoring: To address the dynamic nature of AI systems, where the model’s behavior can change over time due to learning processes, a dynamic scoring system could be implemented. This would involve periodic reassessments of vulnerabilities based on how the AI model evolves. One approach could be to use automated monitoring tools that track changes in the model’s behavior and reevaluate vulnerability scores at defined intervals or when significant changes occur in the model’s input data or learning parameters. Implementing dynamic scoring raises questions about the frequency and triggers for score updates, the stability of vulnerability management programs, and the computational overhead involved in continuously monitoring AI systems.
  • AI-Specific Metrics: Developing AI-specific metrics involves identifying and defining impacts and exploitabilities unique to AI systems. While measuring the impact of AI vulnerabilities on Confidentiality, Integrity, and Availability (CIA) is important, it is also necessary to develop a new set of metrics meaningful for AI software. For instance, metrics could measure susceptibility to data poisoning, resistance to model inversion attacks, and sensitivity to adversarial examples. Another area could be the robustness of the model under distributional shift, where the data’s statistical properties change over time, which is not currently addressed in traditional CVSS metrics. The main challenge is the technical complexity involved in quantifying such metrics, especially in a standardized manner that fits diverse AI applications across different industries.
This area remains open and requires further research within the community to develop a robust and accurate severity scoring system.
@@ II-D Affected Software and Products
Specifies the software products, AI models, or systems that contain or are susceptible to the vulnerability. This may include versions, configurations, or specific components that are affected. This will require additional research to map a reported vulnerability to a unique set of AI products.

### v2

@@ Abstract
In the rapidly evolving field of artificial intelligence (AI), the identification, documentation, and mitigation of vulnerabilities are paramount to ensuring robust and secure systems. This paper discusses minimum elements for AI vulnerability management and the establishment of an Artificial Intelligence Vulnerability Database (AIVD) supported by an AI Bill of Materials (AIBOM) for structured model documentation. It presents standardized formats and protocols for disclosing, analyzing, cataloging, and documenting AI vulnerabilities. It discusses how such an AI incident database must extend beyond the traditional scope of vulnerabilities by focusing on the unique aspects of AI systems. Additionally, this paper highlights challenges and gaps in AI Vulnerability Management, including the need for new severity scores, weakness enumeration systems, and comprehensive mitigation strategies specifically designed to address the multifaceted nature of AI vulnerabilities.
@@ I Introduction
As artificial intelligence (AI) technologies permeate various sectors, including healthcare [1], finance [2], transportation [3], industry[4, 5], and security [6, 7, 8, 9], their transformative impact is undeniable. However, this rapid integration introduces complex vulnerabilities that may compromise the systems it is intended to enhance. With AI becoming a foundational component across various sectors, there is an urgent need to systematically identify, evaluate, and mitigate AI-specific vulnerabilities. Therefore, Vulnerability Management in AI systems is a critical process that encompasses the identification, assessment, public disclosure, and remediation of security vulnerabilities within the systems and components. This practice is essential due to the unique and complex nature of AI software (software that contains AI components), which often includes data-driven decision-making processes that are opaque and challenging to secure.
@@
The establishment of a robust vulnerability management framework for AI systems underscores the need for centralized information sharing, where AI vulnerabilities can be reported, cataloged, and communicated with all stakeholders. This centralization ensures that information regarding new and emerging threats is quickly and widely disseminated, allowing for more rapid responses across the industry. This demands a global AI vulnerability database to serve as a central repository for information on identified vulnerabilities specific to AI systems, providing crucial support for organizations in protecting their AI technologies against security threats. While the National Vulnerability Database (NVD) [10], along with frameworks like Common Vulnerabilities and Exposures (CVE) [11] and Common Weakness Enumeration (CWE) [12], provides a well-established infrastructure for documenting traditional software vulnerabilities, these systems are designed for static, code-based flaws and lack support for the dynamic, data-driven nature of AI. They do not account for vulnerabilities emerging from model behavior, training data quality, or evolving algorithmic interactions.
@@
AI vulnerabilities differ significantly from traditional software vulnerabilities due to the adaptive and often data-driven nature of AI systems. Unlike conventional software, where vulnerabilities typically arise from coding errors or system misconfigurations, AI vulnerabilities often stem from training data, model architecture, or learning algorithms. For instance, adversarial attacks [13], where malicious inputs are crafted to manipulate model outputs, highlight threat vectors unique to AI. Such vulnerabilities can lead AI systems to behave unpredictably or make decisions that benefit attackers. Identifying and mitigating these issues requires analyzing not only the code but also the behavior and data that shape the model, making the process more complex and less deterministic than in traditional systems.
Given these fundamental differences, there is a clear need for a dedicated repository that focuses specifically on AI vulnerabilities. The proposed Artificial Intelligence Vulnerability Database (AIVD) is designed to fulfill this role, functioning similarly to the NVD but tailored to capture the unique characteristics of AI systems. In this paper, we outline the minimum elements (MEs) required for effective AI vulnerability management and propose a framework for establishing AIVD. We introduce standardized formats and protocols for reporting, analyzing, cataloging, and measuring AI vulnerabilities.
Unlike existing databases, AIVD must account for AI-specific issues such as biases in training data and flaws in algorithmic design. We also identify critical gaps in current practices, including the need for new severity scoring mechanisms, weakness classification systems, and mitigation strategies tailored to the evolving behavior of AI systems. These systems, driven by data and algorithmic complexity, require adaptive vulnerability management approaches to account for their dynamic, non-deterministic nature. Section VI further expands on these challenges and outlines the path forward.
The main contributions of this paper are five-fold:
  1. 1.
    AI-specific Common Weakness Enumeration (AI-CWE) Definition: We introduce a structured weakness enumeration system tailored to AI systems, distinct from traditional CWE by focusing on AI-specific failure modes.
  2. 2.
    MEs for AI Vulnerabilities: We define a novel set of MEs necessary for describing AI vulnerabilities across model, data, and deployment layers—supporting consistent reporting and assessment.
  3. 3.
    AI Bill of Materials (AIBoM) Framework: We outline AIBOM, an AI-specific extension to SBOM, capturing not only system components and dependencies but also ethical, environmental, and usage boundaries specific to AI workflows.
  4. 4.
    AI-Aware Severity Scoring: We analyze the gaps in CVSS and propose two directions—dynamic scoring and AI-specific impact metrics—for developing a more suitable scoring system for AI vulnerabilities.
  5. 5.
    Gap Analysis and Future Challenges in AI Vulnerability Management: We identify and formalize the critical challenges that distinguish AI vulnerability management from traditional software security practices.
TABLE I: Details of AI Vulnerability Assessment
@@
1 AI-CVE ID This is a unique identifier assigned to each recorded vulnerability within the AIVD fortracking and referencing.
2 AI Model Details Specific details of the AI model or system affected (e.g., model type, version).
3 Weakness Type Identifies the category or mechanism of the AI system vulnerability. Similar to CWE in NVD database, each vulnerability will be assigned with an AI-CWE.
4 Root Cause Explains factors contributing to the vulnerability, such as inherent model flaws, insufficient robustness safeguards, or data issues.
@@
6 Severity Scores Quantifies the impact of AI vulnerabilities, helping prioritize and guide effective mitigation.
7 Affected Software & Products Lists applications affected by the vulnerability to guide targeted remediation and risk management.
8 Exploitability Describes how the vulnerability can be exploited, including attack complexity, required privileges, and user actions.
@@ II Minimum Elements for Defining AI Vulnerabilities
This section describes the Minimum Elements (MEs) essential for capturing and describing AI vulnerabilities effectively. Each ME represents a critical piece of information that contributes to the comprehensive profiling and understanding of vulnerabilities within AI systems. These elements facilitate structured data collection, enabling systematic tracking, analysis, prioritizing, and mitigation of vulnerabilities. These elements equip stakeholders with the necessary details to address vulnerabilities from identification to resolution. Each ME is strategically defined to encapsulate crucial information regarding the associated vulnerability. This approach not only standardizes vulnerability reporting across various platforms and industries, but also ensures that all stakeholders, from developers and security professionals to regulatory bodies, have access to the same comprehensive data set. Structured data collection enabled by these elements allows organizations to systematically track vulnerabilities from initial identification to final resolution. This process involves analyzing vulnerabilities for impact and exploitability, helping prioritize them for mitigation.
Furthermore, these elements serve as the backbone for creating databases and registries that stakeholders can refer to to ensure consistency and accuracy in vulnerability management. By defining these minimum required data points, the AI vulnerability disclosure process becomes more efficient, transparent, and scalable. This accommodates the dynamic nature of AI technologies and their complex ecosystems. The standardized approach streamlines mitigation and improves AI security by providing clear actionable information for rapid threat response.These minimum elements are described in Table I. Some of these minimum elements require further advancements in foundational knowledge and the creation of new frameworks and catalogs to describe AI vulnerabilities.
@@ II-A Weakness Type
Despite their sophistication, AI systems are uniquely vulnerable to weaknesses rarely seen in traditional software. These vulnerabilities arise from core characteristics of AI, including its reliance on large datasets and complex, often opaque decision-making processes. While resources like CWE catalog traditional software vulnerabilities, there is no equivalent framework for AI. Moreover, CWE cannot fully capture AI-specific weaknesses introduced by data dependence and algorithmic complexity. Because AI performance hinges on data quality and volume, any corruption or bias can severely affect outcomes. The black-box nature of many AI models further complicates vulnerability detection and understanding, hindering the development of robust systems.
To address AI-specific vulnerabilities, we propose an AI-CWE—a catalog of weaknesses enabling a structured approach to their identification. The AI-CWE builds on the AI/ML ATT&CK framework [14], linking each weakness to gaps in mitigation strategies. We define four main classes of weaknesses as follows:
@@
Each weakness is categorized to support detailed analysis, root-cause tracing, and targeted mitigation. Table II illustrates sample AI-CWE entries and their key elements.
TABLE II: Example of a potential AI-CWE entry (inadequate input filtering)
@@ II-B Impact
Describes the repercussions of a vulnerability within the AI system, explaining how it affects functionality, security, and reliability. It covers both direct and indirect effects, including performance degradation, exploitation risks, and broader impacts on users and services.
@@ II-C Severity Scores
The Common Vulnerability Scoring System (CVSS) is a widely used standard for assessing software vulnerability severity. However, in AI systems, unique characteristics such as learning behavior and data dependence reveal gaps in CVSS and similar frameworks. Given the distinctive nature of AI vulnerabilities, a tailored severity assessment approach is needed. Below, we outline these gaps, discuss measurement methods, and propose future directions.
@@ II-C1 CVSS Base Metrics and AI
  • Exploitability Metrics (Attack Vector, Attack Complexity, Privileges Required, User Interaction): Define how a vulnerability can be exploited. AI systems add complexities not captured here, such as subtle data manipulation influencing model behavior.
  • Impact Metrics (Confidentiality, Integrity, Availability): While covering direct impacts, AI systems also face indirect effects—like decision errors or data poisoning—that degrade operational integrity beyond data loss or downtime.
@@ II-C2 Environmental Metrics and AI
These metrics adapt to organizational infrastructures, vital for AI systems whose deployment context shapes vulnerability impact. However, AI’s evolving behavior and data interactions require more flexible metrics.
@@ II-C3 Supplemental Metrics and AI
Safety, Automatable, Recovery, and Value Density metrics add contextual insight. In AI systems, Safety is crucial as decisions affect human well-being, while Automatable is key due to large-scale automated attack risks.
@@ II-C4 Notable Gaps
While CVSS is comprehensive, several key gaps arise when applied to AI: (1) Dynamic Nature of AI: AI systems evolve by learning from new data, altering their vulnerability landscape. Static CVSS scores may not capture this ongoing change. (2) Indirect Impacts: AI vulnerabilities may not directly affect confidentiality, integrity, or availability but can degrade decision quality—impacts not well reflected in current CVSS metrics. (3) Data and Model-Specific Attacks: AI-specific threats like data manipulation and model evasion lack direct analogs in traditional software, demanding new impact and exploitability metrics.
Adapting CVSS for AI thus demands systematic, dynamic assessment of AI-specific risks.
  • Dynamic Scoring: To address AI’s evolving nature, where model behavior changes through learning, a dynamic scoring system is needed. This involves periodic reassessment of vulnerabilities using automated tools to track behavioral shifts and update scores when significant changes occur. Key challenges include defining update frequency, maintaining management stability, and controlling the computational cost of continuous monitoring.
  • AI-Specific Metrics: Developing AI-specific metrics requires defining impacts and exploitabilities unique to AI systems. While measuring vulnerabilities through Confidentiality, Integrity, and Availability (CIA) remains important, new metrics are needed for AI software. These may assess susceptibility to data poisoning, resistance to model inversion, sensitivity to adversarial examples, or robustness under distributional shift—factors not addressed in traditional CVSS metrics. The main challenge lies in quantifying such metrics consistently across diverse AI applications and industries.
This area remains open and requires further research to develop a reliable severity scoring system.
@@ II-D Affected Software and Products
Identifies the software products, AI models, or systems vulnerable to the issue, including relevant versions, configurations, or components. Mapping a reported vulnerability to specific AI products may require further research.